OpenStudy (anonymous):

Solve for n. (8 3/4 + 7 2/5) + 6 3/5 = n + 14 A. 9 3/4 B. 7 4/7 C. 8 3/7 D. 7 3/4

OpenStudy (anonymous):

I will give a medal

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

d i found it online

OpenStudy (anonymous):

Okay lol

OpenStudy (anonymous):

the medal

OpenStudy (anonymous):

thanks

OpenStudy (welshfella):

None of those options are correct There is a mistake in the question somewhere

OpenStudy (anonymous):

I copy and pasted the answer and the question

OpenStudy (welshfella):

Just giving the answer does not teach the questioner anything...

OpenStudy (welshfella):

well none of them are correct if you add up the left side you get 91/4 and 14 + 7 3/4 = 87/4

OpenStudy (welshfella):

correct is 91/4 - 14 = 8 3/4

OpenStudy (welshfella):

maybe C should be 8 3/4 not 8 3/7